a gas has a volume of 255mL, a pressure of 733 mm Hg, a temperature of 548 degrees celcius and a mass of 1.22g. What is the molar mass?
1
Expert's answer
2016-01-29T09:43:54-0500
pV = nRT; n=m/M;
M = mRT/pV

1) If using atmospheres:
Convert mm Hg to atm = 733 mm Hg x 1 atm/760mm Hg = 0.9645 atm
Convert mL to L = 255 mL x 1L/1000mL = 0.255L
Convert C to K = 548 ℃ + 273.15 K = 821.15°K
R = 0.08206 atm/mol*K

M = (1.22g x 0.08206 atm/mol*K x 821.15°K) / (0.9645atm x 0.255L) = 334.25 g/mol

2) If using SI units:
Convert mm Hg to Pa = 733 mm Hg x 101325 Pa/760mm Hg = 97725.294 Pa
Convert mL to m3 = 255 mL x 1m3/106mL = 0.255*10-3 m3
Convert C to K = 548 ℃ + 273.15 K = 821.15°K
R = 8.3145 J/mol*K

M = (1.22g x 8.3145 J/mol*K x 821.15°K) / (97725.294Pa x 0.255*10-3 m3) = 334.25 g/mol